A. Assume that the demand for real money balance (M/P)d is M/P = 0.4Y – 10i. National income is 30,000 and the price level is 100 and the growth rate of nominal money is 2 percent. The real interest rate r is fixed at 3 percent. Also assume that the expected inflation rate equals the rate of nominal money growth.

B. What is the quantity of money in the economy?
Answer #1

The quantity of money in the economy is M

So, we have M/P = 0.4Y – 10i Putting the values of the variables in the equation, we have

M/100=0.4*30000-10*(2+3)=11950

So, we have M=11,950,000